# Best Ergodic Averages via Optimal Graph Filters in Reversible Markov Chains

### v1

@@ IV Filter Design with Bernstein Polynomials
In this sub-section, instead of pursuing the optimal design formulated in ${\bf(OPT_{t})}$, we approach the problem differently. Firstly, the behavior of the polynomial $p$ over the interval $(0,\lambda_{{\mathsf{low}}})$ is not critical. Therefore, let us fix a continuous function $g$ on $[0,2]$ such that $g(\lambda_{x_{*}})=g(0)=1$ and $g(z)=0$ for $2\geq z\geq\lambda_{{\mathsf{low}}}$. We have the freedom to choose the behavior of $g$ on the interval $(0,\lambda_{{\mathsf{low}}})$ as desired. This choice is also part of the design process. Note that Weierstrass approximation theorem [19] states that any continuous function over compact domain can be approximated via polynomials under sup-norm. This theorem is generally proved constructively using Bernstein polynomials (see the proof of [19, Theorem 1.1]). Therefore, we can use Bernstein polynomials to approximate our function $g$ over $[0,2]$. To this end, let us first define Bernstein polynomials.
Definition 9.
Given any function $h$ that is bounded on $[0,1]$, we define its Bernstein polynomial of degree $K$ by
$B_{K}(h;z)\coloneqq\sum_{l=0}^{K}h\left(\frac{l}{K}\right){K\choose l}z^{l}\,(1-z)^{K-l}.$
One can prove that (see [19, Theorem 1.2])
$\|h-B_{K}(h)\|_{\infty}^{[0,1]}\leq\frac{3}{2}\,\omega_{h}\left(\frac{1}{\sqrt{K}}\right),$
where $\omega_{h}$ is the modulus of continuity of $h$. Since our function $g$ is defined over $[0,2]$ and Bernstein polynomials are defined over $[0,1]$, we need to translate this discussion into the interval $[0,2]$. The Bernstein polynomial of degree $K$ of $g$ is defined as
$B_{K}(g;z)\coloneqq\sum_{l=0}^{K}g\left(\frac{2l}{K}\right){K\choose l}\left(\frac{z}{2}\right)^{l}\,\left(1-\frac{z}{2}\right)^{K-l}.$
@@
Additionally, $B_{K}(g;\lambda_{x_{*}})=B_{K}(g;0)=1$. Thus, $B_{K}(g)$ is a strong candidate for our polynomial filter. However, given the above rate of convergence result, to achieve a sharp decrease in the function value of $B_{K}(g)$ over the interval $[\lambda_{{\mathsf{low}}},2]$, it makes sense to select $g$ or adjust its behavior on the interval $(0,\lambda_{{\mathsf{low}}})$ so that $\omega_{g}$ is minimized at $\frac{2}{\sqrt{K}}$. Consequently, we propose the following optimization problem for each time step $t\geq 2$:
@@
Solving this optimization problem for any $t\geq 2$ is generally challenging and impractical to perform at each time step $t\geq 2$. However, as $t$ approaches infinity, we can demonstrate that the triangle function $g_{\Delta}$, which decreases from 1 to zero as we move from 0 to $\lambda_{{\mathsf{low}}}$, is proved to be asymptotically optimal for this problem. Therefore, for each $t\geq 2$, we can select $g$ as $g_{\Delta}$ to design our Bernstein filter.
Asymptotically Optimal Bernstein Filter $B_{K}(g_{\Delta};z)\coloneqq\sum_{l=0}^{K}g_{\Delta}\left(\frac{2l}{K}\right){K\choose l}\left(\frac{z}{2}\right)^{l}\,\left(1-\frac{z}{2}\right)^{K-l}$
@@
where $m\coloneqq 1/\lambda_{{\mathsf{low}}}$. Note that if $2/\sqrt{t-1}\geq\lambda_{{\mathsf{low}}}$, then for any $g\in C[0,2]$ satisfying the constraints in ${\bf(OPT^{B}_{t})}$, we have $\omega_{g}\left(2/\sqrt{t-1}\right)=1$. Hence, any feasible $g\in C[0,2]$ solves ${\bf(OPT^{B}_{t})}$. For this reason, in the following, we assume that $2/\sqrt{t-1}<\lambda_{{\mathsf{low}}}$.
Proposition 1.
If $g\in C[0,2]$ is an optimal solution of ${\bf(OPT^{B}_{t})}$, where $r\coloneqq 2/\sqrt{t-1}<\lambda_{{\mathsf{low}}}$, then
$\omega_{\Delta}(r)\frac{n+\lambda}{n+1+\lambda}\leq\omega_{g}(r)<\omega_{\Delta}(r).$
Here, $\lambda_{{\mathsf{low}}}=nr+\lambda r$ and $\lambda\in(0,1)$. Hence, as $t\rightarrow\infty$, we have $\lambda\rightarrow 0$, $n\rightarrow\infty$, and $\frac{n+\lambda}{n+1+\lambda}\rightarrow 1$. This implies that $g_{\Delta}$ is asymptotically optimal.
Proof.
Let $g\in C[0,2]$ be an optimal solution [^2] If the optimal solution does not exist, we can use an $\varepsilon$-optimal solution for a sufficiently small $\varepsilon>0$. In this case, we should replace $\omega_{\Delta}$ with $\omega_{\Delta}+\varepsilon$ in the second inequality stated in the proposition. of ${\bf(OPT^{B}_{t})}$, where $r<\lambda_{{\mathsf{low}}}$. By [19, Lemma 1.3], we have
11 11 $=\omega_{g}(nr+\lambda r)\leq(n+\lambda+1)\omega_{g}(r)$
$\leq(n+\lambda+1)\omega_{\Delta}(r).$
But note that $1=\omega_{\Delta}(nr+\lambda r)=(n+\lambda)\omega_{\Delta}(r)$. Hence
$(n+\lambda)\omega_{\Delta}(r)\leq(n+\lambda+1)\omega_{g}(r)\leq(n+\lambda+1)\omega_{\Delta}(r).$
Dividing both sides with $(n+\lambda+1)$ establishes the result. ∎
Another justification for why the triangle function is a good candidate is the following. Instead of considering the space $C[0,2]$, we can work with the set of continuously differentiable, except possibly at $\lambda_{{\mathsf{low}}}$, functions on $[0,2]$, denoted as $C^{1}[0,2]$. Since estimating the modulus of continuity exactly can be challenging, we can use the inequality $\omega_{g}(\delta)\leq\|g^{(1)}\|_{\infty}^{[0,2]}\delta$ for all $\delta\geq 0$ to obtain an upper bound on the modulus of continuity. This allows us to state a more manageable optimization problem, which does not depend on time.
@@
Proof.
According to the mean value theorem, for any function $g\in C^{1}[0,2]$ satisfying $g(0)=1$ and $g([\lambda_{{\mathsf{low}}},2])=0$, there exists an $l\in(0,\lambda_{{\mathsf{low}}})$ such that $|g^{(1)}(l)|=1/\lambda_{{\mathsf{low}}}$, and so, $\|g^{(1)}\|_{\infty}^{[0,2]}\geq 1/\lambda_{{\mathsf{low}}}$. Furthermore, note that for the triangle function, $\|g^{(1)}_{\Delta}\|_{\infty}^{[0,2]}=1/\lambda_{{\mathsf{low}}}$. $\square$
V Filter Design with Chebyshev Polynomials
@@
In this section, we establish that the optimal solution of ${\bf(OPT_{t})}$ is a properly normalized Chebyshev polynomial of degree $t-1$, for any $t\geq 2$. Since we have explicit forms of Chebyshev polynomials of any degree, we can easily implement the corresponding filter. Before proving this result, let us briefly discuss Chebyshev polynomials and outline one of their key properties, which will lead us to the optimal solution of ${\bf(OPT_{t})}$. This property can be stated as follows: among all polynomials with a fixed degree and a sup-norm less than $1$ over any interval $[a,b]$, Chebyshev polynomials are the ones that increase as quickly as possible compared to other polynomials outside this interval. Using this property, it becomes clear that Chebyshev polynomials are the optimal solutions of ${\bf(OPT_{t})}$.
Chebyshev polynomials are originally defined over the interval $[-1,1]$. Therefore, we begin by defining these polynomials over this interval and stating their maximum growth property. Then, we translate everything to the interval $[\lambda_{{\mathsf{low}}},2]$ in our problem formulation.
Definition 10.
The Chebyshev polynomial of degree $K$ is defined as follows:
$T_{K}(z)\coloneqq\cos(K\theta),$
where $z=\cos(\theta)$ and $\theta\in[0,\pi]$.
Note that since $\cos(K\theta)$ is a polynomial of $\cos(\theta)$ with degree $K$ [19, Exercise 1.6, p. 43], $T_{K}$ is well-defined. Moreover, since it is a polynomial, once we know its coefficients, we know it as a function for all $z$. The following result states that Chebyshev polynomials grows as rapidly as possible outside of $[-1,1]$.
Theorem 2.
[19, Theorem 1.10, p. 31] If $p\in P_{K}[-1,1]$ and $|z_{0}|\geq 1$, then
$|p^{(l)}(z_{0})|\leq\|p\|_{\infty}^{[-1,1]}\,|T_{K}^{(l)}(z_{0})|,\,\,l=0,1,\ldots,K,$
and in particular, we have
$|p(z_{0})|\leq\|p\|_{\infty}^{[-1,1]}\,|T_{K}(z_{0})|.$
Using above maximal growth condition, we can now prove the main result of this section. Before doing that, let us define the Chebyshev polynomial of degree $K$ over the interval $[\lambda_{{\mathsf{low}}},2]$ using a linear transformation as follows:
$T_{K}^{[\lambda_{{\mathsf{low}}},2]}(z)\coloneqq T_{K}\left(\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\right).$
@@
$p_{K}^{*}(z)\coloneqq\frac{T_{K}^{[\lambda_{{\mathsf{low}}},2]}(z)}{T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)}.$
@@
Theorem 3.
We have $p_{t-1}^{*}\in\mathop{\rm arg\,min}{\bf(OPT_{t})}$ for any $t\geq 2$. Hence, the optimal polynomial filter is given by properly normalized Chebyshev polynomial of degree $t-1$ at each time step $t\geq 2$.
Proof.
Fix any $t\geq 2$. Suppose that there exists $p\in P_{t-1}[\lambda_{{\mathsf{low}}},2]$ such that $p(0)=1$ and $\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}<\|p_{t-1}^{*}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}$. Our goal is to reach a contradiction. Note that Theorem 2 is still true for Chebyshev polynomials over the interval $[\lambda_{{\mathsf{low}}},2]$. Therefore, as $0\notin[\lambda_{{\mathsf{low}}},2]$, we have
$|p(0)|\leq\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|.$
But this implies that
$|p(0)|<\|p_{t-1}^{*}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|=1$
as
$\|p_{t-1}^{*}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}=\frac{1}{\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|}.$
This is a contradiction as we assumed $p(0)=1$. ∎
Note that Chebyshev polynomials over $[-1,1]$ satisfy the following recursive relation [19, p. 55]:
$T_{K+1}(z)=2z\,T_{K}(z)-T_{K-1}(z)\,\,K\geq 1.$
Hence we have
@@
$\alpha_{K}\coloneqq\frac{T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)}{T_{K+1}^{[\lambda_{{\mathsf{low}}},2]}(0)},\,\,\beta_{K}\coloneqq\frac{T_{K-1}^{[\lambda_{{\mathsf{low}}},2]}(0)}{T_{K+1}^{[\lambda_{{\mathsf{low}}},2]}(0)}.$
@@
VI Filter Design with Legendre Polynomials

### v2

@@ 4 Filter Design with Bernstein Polynomials
In this sub-section, instead of pursuing the optimal design formulated in $\bf(OPT_{t})$}, we approach the problem differently. Firstly, the behavior of the ideal low-pass filter over the interval $(0,\lambda_{{\mathsf{low}}})$ is not critical. Therefore, let us fix a continuous function $g$ on $[0,2]$ such that $g(\lambda_{x_{*}})=g(0)=1$ and $g(z)=0$ for $2\geq z\geq\lambda_{{\mathsf{low}}}$; that is, $g$ can be considered as an ideal low-pass filter given $\lambda_{{\mathsf{low}}}$. We have the freedom to choose the behavior of $g$ on the interval $(0,\lambda_{{\mathsf{low}}})$ as desired. This choice is also part of the design process. Note that Weierstrass approximation theorem [15] states that any continuous function over compact domain can be approximated via polynomials under sup-norm. This theorem is generally proved constructively using Bernstein polynomials (see the proof of [15, Theorem 1.1]). Therefore, we can use Bernstein polynomials to approximate our ideal low-pass filter $g$ over $[0,2]$. To this end, let us first define Bernstein polynomials.
Definition 4.1.
Given any function $h$ that is bounded on $[0,1]$, we define its Bernstein polynomial of degree $K$ by $B_{K}(h;z)\triangleq\sum_{l=0}^{K}h\left(\frac{l}{K}\right){K\choose l}z^{l}\,(1-z)^{K-l}.$
One can prove that (see [15, Theorem 1.2]) $\|h-B_{K}(h)\|_{\infty}^{[0,1]}\leq\frac{3}{2}\,\omega_{h}\left(\frac{1}{\sqrt{K}}\right),$ where $\omega_{h}$ is the modulus of continuity of $h$. Since our function $g$ is defined over $[0,2]$ and Bernstein polynomials are defined over $[0,1]$, we need to translate this discussion into the interval $[0,2]$. The Bernstein polynomial of degree $K$ of $g$ is defined as
$B_{K}(g;z)\triangleq\sum_{l=0}^{K}g\left(\frac{2l}{K}\right){K\choose l}\left(\frac{z}{2}\right)^{l}\,\left(1-\frac{z}{2}\right)^{K-l}.$
@@
Additionally, $B_{K}(g;\lambda_{x_{*}})=B_{K}(g;0)=1$. Thus, $B_{K}(g)$ is a strong candidate for our polynomial filter. However, given the above rate of convergence result, to achieve a sharp decrease in the function value of $B_{K}(g)$ over the interval $\lambda_{{\mathsf{low}}},2]$, it makes sense to select $g$ or adjust its behavior on the interval $(0,\lambda_{{\mathsf{low}}})$ so that $\omega_{g}$ is minimized at $\frac{2}{\sqrt{K}}$. Consequently, we propose the following optimization problem for each time step $t\geq 2$:
@@
Solving this optimization problem for any $t\geq 2$ is generally challenging and impractical to perform at each time step $t\geq 2$. However, as $t$ approaches infinity, we can demonstrate that the triangle function $g_{\Delta}$, which decreases from 1 to zero as we move from 0 to $\lambda_{{\mathsf{low}}}$, is proved to be asymptotically optimal for this problem. Therefore, for each $t\geq 2$, we can select our ideal low-pass filter $g$ as $g_{\Delta}$ to design our Bernstein filter.
Asymptotically Optimal Bernstein Filter $B_{K}(g_{\Delta};z)\triangleq\sum_{l=0}^{K}g_{\Delta}\left(\frac{2l}{K}\right){K\choose l}\left(\frac{z}{2}\right)^{l}\,\left(1-\frac{z}{2}\right)^{K-l}$
@@
where $m\triangleq 1/\lambda_{{\mathsf{low}}}$. Note that if $2/\sqrt{t-1}\geq\lambda_{{\mathsf{low}}}$, then for any $g\in C[0,2]$ satisfying the constraints in $\bf(OPT^{B}_{t})$}, we have $\omega_{g}\left(2/\sqrt{t-1}\right)=1$. Hence, any feasible $g\in C[0,2]$ solves $\bf(OPT^{B}_{t})$}. For this reason, in the following, we assume that $2/\sqrt{t-1}<\lambda_{{\mathsf{low}}}$.
Proposition 4.2.
If $g\in C[0,2]$ is an optimal solution of $\bf(OPT^{B}_{t})$}, where $r\triangleq 2/\sqrt{t-1}<\lambda_{{\mathsf{low}}}$, then $\omega_{\Delta}(r)\frac{n+\lambda}{n+1+\lambda}\leq\omega_{g}(r)<\omega_{\Delta}(r).$ Here, $\lambda_{{\mathsf{low}}}=nr+\lambda r$ and $\lambda\in(0,1)$. Hence, as $t\rightarrow\infty$, we have $\lambda\rightarrow 0$, $n\rightarrow\infty$, and $\frac{n+\lambda}{n+1+\lambda}\rightarrow 1$. This implies that $g_{\Delta}$ is asymptotically optimal.
Proof 4.3.
Let $g\in C[0,2]$ be an optimal solution [^2] If the optimal solution does not exist, we can use an $\varepsilon$-optimal solution for a sufficiently small $\varepsilon>0$. In this case, we should replace $\omega_{\Delta}$ with $\omega_{\Delta}+\varepsilon$ in the second inequality stated in the proposition. of $\bf(OPT^{B}_{t})$}, where $r<\lambda_{{\mathsf{low}}}$. By [15, Lemma 1.3], we have
$1=\omega_{g}(nr+\lambda r)\leq(n+\lambda+1)\omega_{g}(r)\leq(n+\lambda+1)\omega_{\Delta}(r).$
But note that $1=\omega_{\Delta}(nr+\lambda r)=(n+\lambda)\omega_{\Delta}(r)$. Hence $(n+\lambda)\omega_{\Delta}(r)\leq(n+\lambda+1)\omega_{g}(r)\leq(n+\lambda+1)\omega_{\Delta}(r).$ Dividing both sides with $(n+\lambda+1)$ establishes the result.
Another justification for why the triangle function is a good candidate for ideal low-pass filter is the following. Instead of considering the space $C[0,2]$, we can work with the set of continuously differentiable, except possibly at $\lambda_{{\mathsf{low}}}$, functions on $[0,2]$, denoted as $C^{1}[0,2]$. Since estimating the modulus of continuity exactly can be challenging, we can use the inequality $\omega_{g}(\delta)\leq\|g^{(1)}\|_{\infty}^{[0,2]}\delta$ for all $\delta\geq 0$ to obtain an upper bound on the modulus of continuity. This allows us to state a more manageable optimization problem, which does not depend on time.
@@
Proof 4.5.
According to the mean value theorem, for any function $g\in C^{1}[0,2]$ satisfying $g(0)=1$ and $g([\lambda_{{\mathsf{low}}},2])=0$, there exists an $l\in(0,\lambda_{{\mathsf{low}}})$ such that $|g^{(1)}(l)|=1/\lambda_{{\mathsf{low}}}$, and so, $\|g^{(1)}\|_{\infty}^{[0,2]}\geq 1/\lambda_{{\mathsf{low}}}$. Furthermore, note that for the triangle function, $\|g^{(1)}_{\Delta}\|_{\infty}^{[0,2]}=1/\lambda_{{\mathsf{low}}}$.
5 Filter Design with Chebyshev Polynomials
@@
In this section, we establish that the optimal solution of $\bf(OPT_{t})$} is a properly normalized Chebyshev polynomial of degree t1t-1, for any $t\geq 2$. Since we have explicit forms of Chebyshev polynomials of any degree, we can easily implement the corresponding filter. Before proving this result, let us briefly discuss Chebyshev polynomials and outline one of their key properties, which will lead us to the optimal solution of $\bf(OPT_{t})$}. This property can be stated as follows: among all polynomials with a fixed degree and a sup-norm less than $1$ over any interval $[a,b]$, Chebyshev polynomials are the ones that increase as quickly as possible compared to other polynomials outside this interval. Using this property, it becomes clear that Chebyshev polynomials are the optimal solutions of $\bf(OPT_{t})$}.
Chebyshev polynomials are originally defined over the interval [1,1][-1,1]. Therefore, we begin by defining these polynomials over this interval and stating their maximum growth property. Then, we translate everything to the interval $\lambda_{{\mathsf{low}}},2]$ in our problem formulation.
Definition 5.1.
The Chebyshev polynomial of degree $K$ is defined as follows: $T_{K}(z)\triangleq\cos(K\theta),$ where $z=\cos(\theta)$ and $\theta\in[0,\pi]$.
Note that since $\cos(K\theta)$ is a polynomial of $\cos(\theta)$ with degree $K$ [15, Exercise 1.6, p. 43], $T_{K}$ is well-defined. Moreover, since it is a polynomial, once we know its coefficients, we know it as a function for all $z$. The following result states that Chebyshev polynomials grows as rapidly as possible outside of [1,1][-1,1].
Theorem 5.2.
[15, Theorem 1.10, p. 31] If $p\in P_{K}[-1,1]$ and $|z_{0}|\geq 1$, then $|p^{(l)}(z_{0})|\leq\|p\|_{\infty}^{[-1,1]}\,|T_{K}^{(l)}(z_{0})|,\,\,l=0,1,\ldots,K,$ and in particular, we have $|p(z_{0})|\leq\|p\|_{\infty}^{[-1,1]}\,|T_{K}(z_{0})|.$
Using above maximal growth condition, we can now prove the main result of this section. Before doing that, let us define the Chebyshev polynomial of degree $K$ over the interval $\lambda_{{\mathsf{low}}},2]$ using a linear transformation as follows:
$T_{K}^{[\lambda_{{\mathsf{low}}},2]}(z)\triangleq T_{K}\left(\frac{2z-2-\lambda_{{\mathsf{low}}}}{2-\lambda_{{\mathsf{low}}}}\right).$
@@
$p_{K}^{*}(z)\triangleq\frac{T_{K}^{[\lambda_{{\mathsf{low}}},2]}(z)}{T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)}.$
@@
Theorem 5.3.
We have $p_{t-1}^{*}\in\mathop{\rm arg\,min}{\bf(OPT_{t})}$ for any $t\geq 2$. Hence, the optimal polynomial filter is given by properly normalized Chebyshev polynomial of degree t1t-1 at each time step $t\geq 2$.
Proof 5.4.
Fix any $t\geq 2$. Suppose that there exists $p\in P_{t-1}[\lambda_{{\mathsf{low}}},2]$ such that $p(0)=1$ and $\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}<\|p_{t-1}^{*}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}$. Our goal is to reach a contradiction. Note that Theorem 5.2 is still true for Chebyshev polynomials over the interval $\lambda_{{\mathsf{low}}},2]$. Therefore, as $0\notin[\lambda_{{\mathsf{low}}},2]$, we have $|p(0)|\leq\|p\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|.$ But this implies that $|p(0)|<\|p_{t-1}^{*}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|=1$ as $\|p_{t-1}^{*}\|_{\infty}^{[\lambda_{{\mathsf{low}}},2]}=\frac{1}{\left|T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)\right|}.$ This is a contradiction as we assumed $p(0)=1$.
Note that Chebyshev polynomials over [1,1][-1,1] satisfy the following recursive relation [15, p. 55]: $T_{K+1}(z)=2z\,T_{K}(z)-T_{K-1}(z)\,\,K\geq 1.$ Hence we have
@@
$\alpha_{K}\triangleq\frac{T_{K}^{[\lambda_{{\mathsf{low}}},2]}(0)}{T_{K+1}^{[\lambda_{{\mathsf{low}}},2]}(0)},\,\,\beta_{K}\triangleq\frac{T_{K-1}^{[\lambda_{{\mathsf{low}}},2]}(0)}{T_{K+1}^{[\lambda_{{\mathsf{low}}},2]}(0)}.$
@@
6 Filter Design with Legendre Polynomials